If three forces of 12N, 15N and 20N respectively that are concurrent are acting at a point and in equilibrium, calculate the angle between them
1
Expert's answer
2021-09-28T11:12:44-0400

Align the 20-N force along the y-axis, then, to keep the body in equilibrium, the horizontal and vertical components of the two other forces must be equal:


"20-12\cos\theta_1-15\cos\theta_2=0,\\\space\\
12\sin\theta_1-15\sin\theta_2=0.\\\space\\
12^2(1-\cos^2\theta_1)=15^2(1-\cos^2\theta_2),\\\space\\
(15\cos\theta_2)^2-(12\cos\theta_1)^2-81=0."

So, we have the upper and lower equations that form quite a pretty system of equations. Solve the system (you may use substitution "12\cos\theta_1=x,\space15\cos\theta_2=y"):


"12\cos\theta_1=7.975,\\
15\cos\theta_2=12.025,\\
\theta_1=48.3°,\\
\theta_2=36.7°."